# Grammar-based Game Description Generation using Large Language Models

### v1

@@ Abstract
To lower the barriers to game design development, automated game design, which generates game designs through computational processes, has been explored. In automated game design, machine learning-based techniques such as evolutionary algorithms have achieved success. Benefiting from the remarkable advancements in deep learning, applications in computer vision and natural language processing have progressed in level generation. However, due to the limited amount of data in game design, the application of deep learning has been insufficient for tasks such as game description generation. To pioneer a new approach for handling limited data in automated game design, we focus on the in-context learning of large language models (LLMs). LLMs can capture the features of a task from a few demonstration examples and apply the capabilities acquired during pre-training. We introduce the grammar of game descriptions, which effectively structures the game design space, into the LLMs’ reasoning process. Grammar helps LLMs capture the characteristics of the complex task of game description generation. Furthermore, we propose a decoding method that iteratively improves the generated output by leveraging the grammar. Our experiments demonstrate that this approach performs well in generating game descriptions.
@@
Large Language Model, Ludii, Game Description Language, Grammar, Game Description Language Generation
@@ I Introduction
Game design and development requires a significant amount of expertise, creating barriers that limit individuals trying to create game experiences in various fields such as education, cultural participation, healthcare, and engineering. To reduce these skill barriers and promote the democratization of the game design process, techniques to reduce the workload and streamline the design process have become increasingly important. Various machine learning (ML) techniques have been successful in automated game design, including MCTS [1], random forests [2], and concept expansion [3]. Deep learning-based ML is often used in the domain of procedural content generation for level design [4, 5, 6]. However, despite the progress of Game Description Language (GDL) frameworks [7, 8, 9, 10] in automated game design, there are few examples of applying deep learning techniques from natural language processing to GDL. This is because deep learning techniques require a large amount of training data, making them less prominent in automated game design, where data is scarce. Further validation of the applicability of modern machine learning techniques in automated game design is needed [11].
In this work, we focus on Large Language Models (LLMs) [12, 13] as a new approach to solving tasks with limited data. LLMs are transformer [14]-based deep learning models with a vast number of parameters, pre-trained on a large corpus of textual data. One of the outstanding abilities of LLMs is their capacity to solve various tasks, such as program synthesis [15, 16], using the capabilities acquired during pre-training without additional training. By providing LLMs with task demonstrations consisting of several input-output pairs $(x^{(i)},y^{(i)})$ and a natural language instruction $x$, LLMs can capture the features of the task, in a process known as in-context learning [17]. In automated game design, by providing LLMs with demonstrations of past high-quality game designs represented in a GDL, LLMs can generate game descriptions $y$ using the extensive knowledge obtained during pre-training.
It is challenging to characterize minor domain-specific concepts and semantics from only a few demonstrations [18], even for LLMs. In this work, we propose augmenting the LLM with a grammar $G$ of a GDL into the game design process, as shown in Fig. 1. A GDL represents the game design space structure well, and its grammar can represent the features of the game design generation task with high data efficiency. The minimum grammar $G[y^{(i)}]\subseteq G$ required to construct $y^{(i)}$ is added to the demonstrations. The generation process consists of two stages: the generation of rules in grammar $G[y]$ and the generation of the game description $y$. Furthermore, we propose decoding methods specialized for each stage, which iteratively improve the output of the LLMs using the grammar $G$.
@@
  • We propose a novel framework for automated game design by combining Large Language Models (LLM) and Game Design Languages (GDL).
  • Our framework introduces the grammar of a GDL into the generation process based on in-context learning with LLMs.
@@
Figure 1: An overview of our framework for grammar-based game description generation. We use the grammar of a Game Description Language (GDL) and Large Language Model (LLM) to generate the game description $y$. First, we generate the grammar $\hat{G}[y]$ required to construct $y$ in the Rule Decoding stage, and then generate $\hat{y}$ based on $\hat{G}[y]$ in the Game Description Decoding stage. The core of our framework consists of iteratively decoding to improve the generation, and employing LLMs and the explicit GDL grammar.
@@ II-A Game Description Language
Game Description Language (GDL) is a domain-specific language for games. GDL [9] was introduced in 2005 and has become the standard in General Game Playing, where artificial agents are developed to play a wide variety of games. After GDL, various types of GDL have been developed. VGDL [8] is a language that represents both the levels and rules of 2D sprite-based games. RBG [10] models complex board games by combining low-level and high-level languages. Ludii [7] is a game system developed based on the ludemic approach, which decomposes games into conceptual units of game-related information. Ludii models more than 1,000 traditional games, including board games, card games, dice games, and tile games. The Ludii system generates a complete game description in the Extended Backus-Naur Form (EBNF) style. Diverse analyses of games using Ludii, especially focusing on board games, have been conducted [19, 20, 21, 22, 23]. For example, [19] calculates the distance between board games using concept values extracted from Ludii. Matthew et al. [20] proposes a framework for automatically generating board game manuals using Ludii. In our work, we use Ludii as our GDL. Furthermore, unlike these works, we add a new perspective to the analysis by focusing on game description generation using grammar.
@@ II-B Automated Game Design
Automated game design [24, 25] is one of the core themes in the field of game AI. As AI technology rapidly advances, various aspects of how AI can be utilized in automated game design have been discussed, including the design process [26], design patterns [27], and creative machine learning [28]. Several studies [29, 30] propose methods to generate a game by integrating multiple elements such as visuals, audio, narrative, levels, rules, and gameplay. Generating games using new representations such as answer set programming [31] and game graphs [3] is also explored.
Automated game design for generating game designs in the VGDL [8] format has been explored [32, 33, 1, 2]. Thorbjørn et al. [32] explores the approach to generate VGDL for arcade games using evolutionary algorithms. GVG-RG [1] proposes a framework for generating appropriate game rules for given game levels. Cicero [33] is a mixed-initiative tool that assists in prototyping 2D sprite-based games using VGDL. Thomas et al. [2] aims to acquire a fitness function that guides game design generation using adversarial random forest classifiers. Our approach shares the use of VGDL with these works but differs in exploring new possibilities for automated game design using LLMs.
@@ II-C Large Language Models in Games
Since the advent of ChatGPT [34] in late 2022, Large Language Models (LLMs) have attracted significant attention, and various ways of using LLMs in the game domain are being explored [11]. The goal of procedural content generation (PCG) is to generate content used in digital games, such as rules and levels, and LLMs are frequently tested in this field [4, 5, 6, 35, 36]. Several works [4, 5, 6] use LLMs to generate 2D tile-based game levels. Another work [35] trains a GPT model to generate quests for role-playing games. LLMGG [36] generates both the rules and levels of games represented in VGDL using LLMs. In LLMGG, LLMs take a part of the VGDL [8] representation or examples of other games as prompts, which generate a complete VGDL-based game in one step. The authors discuss that incorrect game levels and rules not included in the VGDL grammar are generated. In contrast, our approach involves multi-step generation, and the necessary grammar rules to build a game are generated as intermediate representations in the middle steps. Additionally, grammar rules and games are generated conditioned on Ludii [7]’s grammar, preventing incorrect syntax and improving consistency.
@@ II-D Program Synthesis
In program synthesis, the task of generating programs from natural language is called semantic parsing. Semantic parsing has benefited from the advancements in LLMs. Several efforts [15, 16] have already explored generating code in general-purpose programming languages such as Python using LLMs. To improve the accuracy of the generated programs, constrained decoding [37, 38, 39] has also been studied. In domain-specific language (DSL) generation, a prompting method [18] has been proposed that introduces grammar as an intermediate product while LLMs iteratively reason to solve tasks (also known as a chain-of-thought [40]). However, in [18], the evaluation targets simple and short DSLs such as SMCalFlow [41] and GeoQuery [42], and the use of LLMs in the approach is limited. We aim to make more effective use of LLMs in our approach and evaluate it with complex and lengthy DSLs that represent game design, such as Ludii [7].
III Background
In this section, we define our problem setting. We first describe the generation of games written in Game Description Language (GDL). We next review in-context learning in game generation.
@@ III-A Game Description Generation
Let $\Sigma^{\ast}$ be the set of all finite strings composed of the alphabet set $\Sigma$. We center our work on a language $D\in\Sigma^{\ast}$ that cam be used to describe games, that is, a GDL. Thus, our goal is to input a natural language query $x$ that describes the content and rules of a game, and generate a corresponding game description $y\in D$. In particular, we use Ludii [7] as our GDL. An example of a Ludii game description generation is shown in Fig. 2.
Figure 2: An example of Ludii game description for the game “Tic-Tac-Toe”.
Ludii is a game system that models over 1,000 traditional games, including board games, card games, dice games, and tile games. Ludii can generate complete game descriptions in Extended Backus-Naur Form (EBNF) syntax. Ludii’s game descriptions include domain-specific abstract concepts, known as ludemes, and their corresponding semantics. We assume full access to the Ludii grammar, denoted as $G$, which has been developed to model a diversity of different games. Let $L(G)$ be the language generated by grammar $G$. Only those game descriptions that are semantically valid become playable games, which we denote as $D$. In other words, we define compilable game descriptions as $L(G)$ and playable game descriptions as $D$.
@@ III-B In-Context Learning
Providing a pretrained Large Language Model (LLM) with a few examples related to a specific task is an efficient way of obtaining more precise and accurate results, and is called in-context learning. This approach does not require additional training or fine-tuning; instead, it relies on the LLM’s ability to identify and apply patterns from the provided examples. In in-context learning, LLM is conditioned on $N$ demonstrations $(x^{(i)},y^{(i)})_{i=1}^{N}$ followed by a test example query $x$, and generates $y$ as $P_{\mathrm{LLM}}(y|(x^{(i)},y^{(i)})_{i=1}^{N},x)$. Recent studies [43, 44] have reported that the few-shot performance on complex reasoning tasks can be improved by inserting intermediate reasoning steps between $x^{(i)}$ and $y^{(i)}$ in the demonstrations.
The effectiveness of in-context learning depends on how effectively the solution to a task can be conveyed through demonstrations. Intuitively, providing more demonstrations to the LLM seems to be beneficial. However, the context length, which is the maximum length that the LLM can capture, is determined during pre-training. Therefore, it is not possible to input a number of demonstrations that exceed this context length. For example, when applying Llama3 [12], one of the leading open-source LLMs, to Ludii’s game descriptions, the context length can easily be exceeded. In particular, the average token length of Ludii’s game descriptions using the Llama3-8B-Instruct tokenizer is 2,458. The context length of Llama3-8B-Instruct is 8,192, which limits the number of Ludii demonstrations that can be input to a few at most. This limitation when using complex examples such as those described in Ludii is common to many LLMs. It is challenging to capture the structure of the wide output space of Ludii’s grammar $G$ from a few demonstrations.
@@ IV Methodology
Figure 4: An example of grammar-based game description generation result. We generate game descriptions in two stages: first we generate the required grammar, and then generate the game description based on the grammar. The prompt includes demonstrations for in-context learning and the test example query $x$, and the demonstration contains the grammar $G[y^{(i)}]$. In the first stage, the minimal grammar $\hat{G}[y]$ that composes $y$ is generated. In the second stage, game description $\hat{y}$ is generated based on the generated $\hat{G}[y]$.
We introduce grammar into intermediate reasoning steps as an extension of in-context learning. A simple approach to game description generation using in-context learning is to provide several examples of a query $x^{(i)}$ and a game description $y^{(i)}$, along with a test query $x$, to a LLM and have the LLM generate a game description $y$. However, this simple approach does not produce decent results. We believe this is because texts related to Ludii are scarcely included in the training corpus for LLMs, and the model has no information about the grammatical structure of Ludii.
In contrast, Backus-Naur Form (BNF) grammars are commonly used to define the structure of languages and are commonly found in the training data of LLM. Inspired by Bailin et al. [18], we incorporate knowledge of BNF grammars in our generation process. Our approach aims to help the LLM capture the structure of the Game Design Language (GDL) by adding the minimal BNF grammar $G[y^{(i)}]$ necessary to compose $y^{(i)}$ in the demonstrations. An example of $G[y]$ for Tic-Tac-Toe is shown in Fig. 3.
Our generation process consists of two stages: first, we input the test example query $x$ into the LLM to generate the minimal grammar $\hat{G}[y]$ necessary to compose $y$. Next, based on $\hat{G}[y]$, the LLM generates the game description $\hat{y}$. To improve the generation performance, we propose methods to decode BNF grammar rules $\hat{G}[y]$ and game descriptions $\hat{y}$ iteratively using the fully accessible Ludii grammar $G$ as our GDL. Afterwards, the decoder evaluates whether $\hat{G}[y]$ (or $\hat{y}$) conforms to the Ludii grammar $G$ (or $\hat{G}[y]$). In each stage, this generation and evaluation process is repeated. In this section, we first explain the two-stage generation in-depth, and then describe the decoding process at each stage in Subsection IV-B.

### v2

@@ Abstract
Game Description Language (GDL) provides a standardized way to express diverse games in a machine-readable format, enabling automated game simulation, and evaluation. While previous research has explored game description generation using search-based methods, generating GDL descriptions from natural language remains a challenging task. This paper presents a novel framework that leverages Large Language Models (LLMs) to generate grammatically accurate game descriptions from natural language. Our approach consists of two stages: first, we gradually generate a minimal grammar based on GDL specifications; second, we iteratively improve the game description through grammar-guided generation. Our framework employs a specialized parser that identifies valid subsequences and candidate symbols from LLM responses, enabling gradual refinement of the output to ensure grammatical correctness. Experimental results demonstrate that our iterative improvement approach significantly outperforms baseline methods that directly use LLM outputs. Our code is available at https://github.com/tsunehiko/ggdg
@@
Large Language Model, Ludii, Game Description Language, Grammar, Game Description Generation
@@ I Introduction
A Game Description Language (GDL) [1, 2, 3, 4, 5] is a domain-specific language that expresses a wide range of games in a unified notation. For example, Ludii GDL [5] models over 1,000 games, primarily board games, as shown in Fig. 1. Game descriptions represented in GDLs are highly machine-readable, making it easy to simulate gameplay using dedicated game engines. Given the amenability of GDLs for automatic game evaluation, they have been extensively used in research on automated game design. In particular, search-based methods such as evolutionary algorithms [4], MCTS [6, 7], and random forests [8] have proven successful in generating game descriptions. Most research primarily focused on mutating existing games based on fitness functions to generate novel games. However, the task of generating game descriptions from natural language texts has not yet been sufficiently explored, and has the potential to lower the bar of entry to game design to non-specialists. In this research, we use Large Language Models (LLMs) [9, 10], which excel at understanding textual context, to generate game descriptions from natural language text in a two-stage process to enforce grammatical correctness.
Figure 1: An example of Ludii game description for the game “Tic-Tac-Toe”. $x$ is text that explains games in natural language. $y$ is a game description in Ludii GDL, a Game Description Language.
LLMs are language models with an enormous number of parameters, pre-trained on vast amounts of text data. These models possess the ability to solve various tasks without additional training [11, 12], and this ability can be elicited by including the task context in the prompt, a technique known as In-Context Learning (ICL) [13]. Hu et al[14] applied this capability to game description generation by incorporating explanations for GDL notations and examples of game descriptions in the prompt context. Their results have shown that more accurate game descriptions can be generated by appropriately refining the prompt context. However, LLMs may still generate grammatically incorrect game descriptions. Such grammatically inaccurate game descriptions cannot be correctly parsed and loaded by game engines, making it difficult to evaluate them through gameplay simulation. To generate higher-quality games, it is important first to ensure that LLMs can produce grammatically correct GDL game descriptions.
This paper presents a method for LLMs to generate more grammatically accurate game descriptions. We propose an approach to iteratively improve LLMs’ initial responses using the GDL grammar. Our generation framework consists of two stages: (i) generating the minimal grammar required for game descriptions, and (ii) iteratively improving the game description based on this minimal grammar. First, we use LLMs to generate the minimal grammar required to produce the game description, making use of the GDL grammar. Next, a parser based on the minimal grammar determines grammatically valid subsequences and a set of candidate symbols that could follow these subsequences from the LLMs’ responses. LLMs then re-infer the missing parts based on these subsequences and candidate symbols, gradually generating grammatically accurate game descriptions. Experimental results demonstrate that our framework is more effective in generating game descriptions compared to a baseline that directly uses the LLMs’ initial responses as an output.
@@
  • We propose a framework for generating game descriptions from natural language text by using LLMs and GDLs.
  • Our framework incorporates GDL grammar into the generation process and iteratively improves the grammatical correctness of the LLM’s output.
@@
Figure 2: An overview of our framework for grammar-based game description generation. We generate the game description $y$ from a natural language query $x$ using Large Language Models (LLMs). First, we generate the grammar $\hat{G}[y]$ required to construct $y$ in the Rule Decoding stage, and then generate $\hat{y}$ based on $\hat{G}[y]$ in the Game Description Decoding stage. The core of our framework is iteratively decoding by leveraging the grammar of game description languages (GDLs) to improve the initial response from LLMs. We make use of Ludii GDL as our GDL, which can model a larger variety of games and is a context-free grammar.
@@ II-A Game Description Language
A Game Description Language (GDL) is a domain-specific language for games. GGP-GDL [1] was introduced in 2005 and has become the standard in General Game Playing, where artificial agents are developed to play a wide variety of games. After GGP-GDL, various types of GDLs have been developed. VGDL [2] is a language that represents both the levels and rules of 2D sprite-based games, and it models 195 games. RBG [3] models complex board games by combining low-level and high-level languages. Ludii [5] is a game system developed based on the ludemic approach, which decomposes games into conceptual units of game-related information. Ludii models more than 1,000 traditional games, including board games, card games, dice games, and tile games. Ludii’s grammar is a Context-Free Grammar (CFG) in Extended Backus-Naur Form (EBNF) style [15]. Due to its ability to generalize and model more games, as well as its capacity to express complete game descriptions in CFG, we use Ludii GDL as our GDL. Note that our approach can be applied to other GDLs that follow CFG, e.g., the rule section of VGDL.
Diverse analyses of games using Ludii, especially focusing on board games, have been conducted [16, 17, 18, 19, 20]. For example, the distance between board games using concept values extracted from Ludii has been proposed in [16]. Stephenson et al. [17] have presented a framework for automatically generating board game manuals using Ludii. Unlike these works, we add a new perspective to the analysis by focusing on game description generation using grammar.
@@ II-B Automated Game Design
Automated game design [21, 22] is one of the core themes in the field of game AI. As AI technology rapidly advances, various aspects of how AI can be utilized in automated game design have been discussed, including the design process [23], design patterns [24], and creative machine learning [25]. Deep learning-based ML is often used in the domain of procedural content generation for level design [26, 27, 28, 29, 30, 31]. Several studies [32, 33] propose methods to generate a game by integrating multiple elements such as visuals, audio, narrative, levels, rules, and gameplay. Generating games using new representations such as answer set programming [34] and game graphs [35] has also been explored. Word2World [36] uses LLMs to design from stories to playable games procedurally. These approaches design games without including GDLs.
Automated game design for generating game designs in the GDL format has been explored [4, 6, 37, 7, 8]. The Ludi system [4], which was the precursor to the ludemic approach used in later Ludii [5], employed evolutionary game design and generated the commercially viable game Yavalath. Thorbjørn et al. [6] explored the approach to generate VGDL [2] for arcade games using evolutionary algorithms. GVG-RG [7] proposes a framework for generating appropriate game rules for given game levels. Cicero [37] is a mixed-initiative tool that assists in prototyping 2D sprite-based games using VGDL. Thomas et al. [8] aimed to acquire a fitness function that guides game design generation using adversarial random forest classifiers. These approaches generate novel games based on existing games, but our approach differs in that it generates game descriptions from natural language text.
@@ II-C Large Language Models in Games
Since the advent of ChatGPT [38] in late 2022, Large Language Models (LLMs) have attracted significant attention, and various ways of using LLMs in automated game design have been explored [39, 40]. Several studies [26, 27] have fine-tuned GPT-2 [41] models to generate 2D tile-based game levels. Prompt-based approaches for LLMs have also been proposed for level generation [28]. Other researchers have trained GPT models to generate quests for role-playing games [42]. Dreamcraft [43] is a method that uses LLMs to generate 3D game objects for Minecraft from text prompts. In addition to generation, researchers have focused on evaluating LLMs. Studies have assessed how well prompt-based level generation methods replicate and generalize [44], and have evaluated the capabilities of LLMs in mixed-initiative game design [45].
GAVEL [46] uses LLMs fine-tuned on Ludii’s game descriptions as mutation operators in evolutionary search. While GAVEL aims to generate games with high novelty, our goal is to generate game descriptions that align with natural language text. LLMGG [14] generates both the rules and levels of games represented in VGDL using LLMs. In LLMGG, LLMs take a part of the VGDL [2] representation or examples of other games as prompts, which generate a complete VGDL-based game in one step. The authors discuss that incorrect game levels and rules not included in the VGDL grammar are generated. In contrast, our approach involves multi-step generation, and the necessary grammar rules to build a game description are generated as intermediate representations in the middle steps. Moreover, since the game description is iteratively generated based on these grammatical rules, this approach prevents inaccurate syntax and improves consistency.
@@ II-D Program Synthesis
In program synthesis, the task of generating programs from natural language is called semantic parsing. Semantic parsing has benefited from advancements in LLMs. Several efforts [11, 12] have already explored generating code in general-purpose programming languages such as Python using LLMs. To improve the accuracy of the generated programs, constrained decoding [47, 48, 49] has also been studied. Grammar-constrained decoding [50, 51, 52] restricts the output space of LLMs to a space described by a grammar, enabling the generation of structured outputs like programs. Our framework includes decoding for grammar generation, in addition to grammar-constrained decoding for game descriptions.
Evolutionary algorithms that use LLMs as evolutionary operators [53, 54, 55, 56] are also gaining attention in program synthesis. Quality-diversity algorithms utilizing LLMs have been proposed for controllable program synthesis, such as neural architecture search [57]. These studies open up the possibility that using LLMs can also improve GDL generation.
In domain-specific language (DSL) generation, a prompting method [58] has been proposed that introduces grammar as an intermediate product while LLMs iteratively reason to solve tasks, also known as a chain-of-thought reasoning [59]. However, in [58], the evaluation targets simple and short DSLs such as SMCalFlow [60] and GeoQuery [61], and the use of LLMs in the approach is limited. We aim to make more effective use of LLMs in our approach and evaluate it with complex and lengthy DSLs that represent game design, such as Ludii [5].
III Problem Setting
TABLE I: Glossary of notations used in this paper.
Notation Description
$G$ GDL grammar (Ludii grammar in this paper)
$L(G)$ Set of game descriptions generated by $G$
$x$ Texts explaining the rules of a game
$y$ A game description (ground truth)
$G[y]$ Minimal grammar extracted from $G$ containing only the grammar rules necessary to generate game description $y$
$(x^{(i)},y^{(i)})_{I=1}^{N}$ Demonstrations of game description generation to be included in LLMs’ prompts
$\hat{G}[y]_{\mathrm{valid}}$ Set of grammar rules in the predicted $\hat{G}[y]$ that are included in $G$
$N_{U}$ Set of non-terminal symbols defined in $G[y]$ but not defined in $\hat{G}[y]_{\mathrm{valid}}$
$G_{N_{U}}$ Set of grammar rules in $G$ defining $N_{U}$
$G[y]_{N_{U}}$ Minimal set of rules extracted from $G_{N_{U}}$ necessary to generate $y$
$\hat{y}_{\mathrm{valid}}$ Part of $\hat{y}$ that conforms to $\hat{G}[y]$
$\Omega_{\mathrm{candidate}}$ Set of candidate terminal symbols to follow $\hat{y}_{\mathrm{valid}}$
$\omega$ Terminal symbol selected from $\Omega_{\mathrm{candidate}}$ by generators such as LLMs
$\hat{\cdot}$ All hat notations indicate symbols are predictions
$\cdot^{\prime}$ All prime notations indicate that the symbols are in the LLM’s output space.
Among these, the one that maximizes the generation probability receives the hat notation.
In this section, we define our problem setting. We first describe the generation of games written in GDLs. We next review in-context learning in the game description generation. Notations used in this paper are summarized in Tab. I.
@@ III-A Game Description Generation
Let $G$ be the GDL grammar, and let $L(G)$ represent the set of game descriptions generated by $G$. We call the task of inputting a natural language query $x$ that describes the content and rules of a game, and generating a corresponding game description $y\in L(G)$, game description generation. Our goal is to make the generated game description $\hat{y}$ as close as possible to the ground truth $y$.
In this paper, we use Ludii GDL as our GDL as it is one of the main references in GDL research. Additionally, Ludii’s grammar is a CFG in EBNF style, which allows us to generate complete game descriptions based on a CFG. An example of Ludii game description generation is shown in Fig.  1. From this point forward, unless otherwise specified, $G$ represents Ludii’s grammar.
@@ III-B In-Context Learning
In-Context Learning (ICL) [13] is an efficient method that provides pretrained LLMs with a few task-specific examples to obtain more precise and accurate results. This approach does not require additional training or fine-tuning; instead, it relies on the LLM’s ability to identify and apply patterns from the provided examples. In ICL, LLM is conditioned on $N$ demonstrations $(x^{(i)},y^{(i)})_{i=1}^{N}$ followed by a test example query $x$, and generates $y$ as $P_{\mathrm{LLM}}(y|(x^{(i)},y^{(i)})_{i=1}^{N},x)$. Recent studies [62, 63] have reported that the few-shot performance on complex reasoning tasks can be improved by inserting intermediate reasoning steps between $x^{(i)}$ and $y^{(i)}$ in the demonstrations.
The effectiveness of ICL depends on how effectively the solution to a task can be conveyed through demonstrations. Intuitively, providing more demonstrations to the LLM seems to be beneficial. However, the context length, which is the maximum length that the LLM can capture, is determined during pre-training. Therefore, it is not possible to input a number of demonstrations that exceed this context length. For example, when applying Llama3 [9], one of the leading open-source LLMs, to Ludii’s game descriptions, the context length can easily be exceeded. In particular, the average token length of Ludii’s game descriptions using the Llama3-8B-Instruct tokenizer is 2,458. The context length of Llama3-8B-Instruct is 8,192, which limits the number of Ludii demonstrations that can be input to a few at most. This limitation when using complex examples such as those described in Ludii is common to many LLMs. Game descriptions generated based on such limited context lack grammatical accuracy and cannot make the games functional and will be discussed in Sec. VI.
@@ IV Methodology
Figure 4: An example of grammar-based game description generation result. We generate game descriptions in two stages: first we generate the required grammar, and then generate the game description based on the grammar. The prompt includes demonstrations for in-context learning and the test example query $x$, and the demonstration contains the grammar $G[y^{(i)}]$. In the first stage, the minimal grammar $\hat{G}[y]$ that composes $y$ is generated. In the second stage, game description $\hat{y}$ is generated based on the generated $\hat{G}[y]$. The red arrow indicates that $\hat{G}[y]$ generated in the first stage is included in the prompt for the second stage.
In this section, we explain our approach to iteratively improve the initial responses of the LLM. To evaluate the game description and improve its grammatical accuracy, we introduce the minimum grammar $G[y]\subseteq G$ required to construct $y$. This minimum grammar $G[y]$ is expected to provide more context to LLMs when included in the prompt, thereby enhancing In-Context Learning (ICL). As shown in Fig. 2, our generation process consists of two stages: First, we input the query $x$ of the test example into an LLM to generate the minimum grammar $\hat{G}[y]$ required to construct $y$. $\hat{G}[y]$ is then evaluated, and the parts that conform to the Ludii grammar $G$ are extracted. The LLM then generates the missing rules. Next, based on $\hat{G}[y]$, the LLM generates the game description $\hat{y}$. $\hat{y}$ is evaluated, and the parts that conform to $\hat{G}[y]$ are extracted. The LLM then infers the rest. Evaluation and LLM generation are repeated at each stage. In this section, we first introduce the minimum grammar $G[y]$, then explain the two-stage generation, and finally describe the decoding process for each stage in detail.